# Einstein Telescope: Detection of Binary Black Hole Gravitational Wave Signals Using Deep Learning

### v3

@@ 3.1 ResNet training
Table 1 displays the total number of TSDCD samples utilized for ResNet model training, testing, and validation. All $F_{low}$ values and $MSNR$ ranges are evenly represented in the randomly chosen testing and validation sources, to ensure an unbiased evaluation of the model’s performance. The three sub-detector combined spectrograms consist of three layers, where each layer belongs to one sud-detector. Hence, the input layer of ResNet was adjusted to a shape size of 365 × 42 × 3, which describes the width, height, and number of layers of the input image, respectively. A batch size of 256, learning rate of 0.0001, and RMSprop optimizer (Vijendra Babu et al., 2020; Xu et al., 2021; Elshamy et al., 2023) were used for training the ResNet model for 200 epochs.
@@ 4 TSDCD evaluation
To compare the TSDCD results with those obtained in our prior investigations in WTM1, we generated a TSDCD with an $F_{low}$ of 30 Hz only with all 5 $MSNR$ ranges. This dataset consists of 4000 sources, created using identical parameters to the testing dataset in WTM1. In Figure 5, the top panel displays the accuracy of the detection acquired for the TSDCD, alongside the accuracy obtained by SSDD in WTM1. The plot clearly demonstrates a significant improvement across all five $MSNR$ ranges, particularly at lower $MSNR$ values. The detection accuracy has improved from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$, and $98.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and greater than 8, respectively. This shows a significant improvement of approximately $30.83\%$, $39.00\%$, $17.75\%$, $5.81\%$ and $1.52\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and greater than 8, respectively. It is essential to note that achieving $100\%$ accuracy for sources with $MSNR$ greater than 8 means the successful detection of all sources within this specific range.
@@
5 Evaluation on ET’s mock data
As previously done in WTM1, three sub-detectors combined mock data was generated to assess the efficiency of utilizing the TSDCD for near-real-time detection and to compare it with previously obtained results from SSDD. Five hours and 50 seconds of time-series data, encompassing all $MSNR$ ranges, were generated for each $F_{low}$ (5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz) for each detector. At every five-second interval in this time-series data, a signal is injected at a random position, resulting in a total of 3,610 injected sources, with each $MSNR$ range containing 722 sources. The total duration of the data for each detector is 25 hours, 4 minutes, and 16 seconds. To compare with our previously obtained results, we utilize the data with an $F_{low}$ of 30 Hz.
Figure 8: Inferencing on ET’s mock data
@@ 5 Evaluation on ET’s mock data
Figure 9: Top: A comparison of FAR values between SSDD (from WTM1) and TSDCD for sources with $F_{low}$ set at 30 Hz. Bottom: FAR values exclusively from TSDCD for sources with $F_{low}$ at 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz.
@@
FPR is equivalent to the False Alarm Rate (FAR), a common measure in GW detection. FAR quantifies how often a GW detector produces a false detection (a false positive) when there is no actual GW signal present (Morrás et al., 2023). This rate is calculated per unit of data length. Therefore, the term FAR will be used instead.
FAR was computed for each $F_{low}$ for the entire mock data. Figure 9 depicts FAR as a function of $MSNR$, where in the top panel, TSDCD results are compared to SSDD results from WTM1. The general trend is consistent with the previous results in WTM1, where FAR decreases as $MSNR$ increases. As shown in the figure, FAR values have significantly improved, especially for sources with lower $MSNR$, transitioning from 0.712, 0.535, 0.050, 0.014, and 0.001 to 0.526, 0.296, 0.006, 0.004, and 0.0 for $MSNR$ ranges 4-5, 5-6, 6-7, 7-8, and >8, respectively. In the bottom panel, FAR values are depicted for each $F_{low}$. The overall performance is nearly identical across different $F_{low}$ settings, except for the case of $F_{low}$ at 5 Hz. This discrepancy aligns with the accuracy results presented earlier.
@@
Similarly, in terms of chirp mass $M$, the right panel in Figure 11 demonstrates a significant decrease in the total number of undetected sources in TSDCD, particularly among sources with chirp masses $M$ in the range of 15-20 $M_{\odot}$. In TSDCD alone, count of undetected sources of different $F_{low}$ varies among all chirp mass $M$ values. For $F_{low}$ at 10 Hz, the minimum count observed is 240, 98, 46,55 and 34 in the chirp mass $M$ ranges of 15-20, 20-25, 25-30,30-35 and 35-40 $M_{\odot}$, respectively. For the chirp mass $M$ ranges 40-45 and 45-50, the minimum count observed are 14 and 7 for $F_{low}$ of 20 Hz and 15 Hz, respectively. Regarding the $D_{L}$, the right panel in Figure 12 shows a substantial reduction in the count of undetected sources in TSDCD, especially for sources with $D_{L}$ in the range of 10-30 Gpc. In the left panel, no significant variation is observed across different $F_{low}$ settings. In summary, both SSDD and TSDCD exhibit a higher detection rate for sources with higher $MSNR$, chirp mass, and shorter $D_{L}$, as indicated by smaller FAR values. Notably, TSDCD demonstrates a significant improvement over SSDD. It’s worth noting that the same source may exhibit different $MSNR$ values on each sub-detector, yet TSDCD consistently outperforms SSDD across these variations.
@@
In terms of computational efficiency, we concurrently processed a total of 25 hours, four minutes, and 16 seconds of data from each detector in parallel. Utilizing the same configuration as in WTM1 (Core i7 MacBook Pro with 16 GB memory, 2667 MHz DDR4, and 2.6 GHz processor), the entire data scan was completed in 22 minutes, averaging 4.7 minutes for each hour. The processing time remains consistent with that achieved on SSDD. The key difference lies in simultaneously reading data from all three sub-detectors, ensuring the computational cost remains unchanged. Hence, utilizing TSDCD is suitable for near-real-time detection and could be further enhanced with a more powerful setup.
@@ 6 Conclusion
In this study, we conducted a comparative analysis of the efficiency in detecting BBH sources using data from all three sub-detectors of ET simultaneously versus using data from a single sub-detector. The comparison was specifically focused on cases with $F_{\text{low}}$ set to 30 Hz. Furthermore, we explored TSDCD’s performance across five different $F_{\text{low}}$ settings (5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz), employing the same $MSNR$ ranges as in our previous study (4-5, 5-6, 6-7, 7-8, and >8). Utilizing the ResNet model, which exhibited superior performance in our previous study, TSDCD demonstrated a significant enhancement in detection accuracy compared to SSDD. The accuracy improved from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$, and $98.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ ranges of 4-5, 5-6, 6-7, 7-8, and >8, respectively. The results indicate a substantial accuracy improvement for lower $MSNR$ ranges: 4-5, 5-6, and 6-7, with gains of $18.5\%$, $24.5\%$, and $13\%$ respectively. Additionally, there is a significant improvement of $5.5\%$ and $1.5\%$ for higher $MSNR$ ranges: 7-8 and >8. In qualitative evaluation, the ResNet model detected sources at 86.601 Gpc with an average $MSNR$ of 3.9 (averaged over the three detectors) and a chirp mass of 13.632 at 5 Hz. The complete scan of 25 hours of data (for each sub-detector) took 22 minutes, which equates to 4.7 minutes for each hour. This makes the use of TSDCD appropriate for near-real-time detection and can be greatly improved using a more powerful setup. Future work will focus on including binary neutron stars (BNS) and verifying the performance on real data from LIGO, VIRGO, and KAGRA.
@@
This research was funded in part by National Science Centre, Poland (UMO-2023/49/N/ST9/01287). Additionally, This work has been supported in part by the European Union’s Horizon 2020 research and innovation programme under grant agreement No 952480 (DarkWave project), and from the International Research Agenda Programme AstroCeNT (MAB/2018/7) funded by the Foundation for Polish Science (FNP) from the European Regional Development Fund. The authors thank Prof. Piotr Gawron for his helpful discussions and reviews.

### v6

@@ 3.1 ResNet training
Table 1 displays the total number of TSDCD samples utilized for ResNet model training, testing, and validation. All $F_{low}$ values and $MSNR$ ranges are evenly represented in the randomly chosen testing and validation sources, to ensure an unbiased evaluation of the model’s performance. The three sub-detector combined spectrograms consist of three layers, where each layer belongs to one sub-detector. Hence, the input layer of ResNet was adjusted to a shape size of 365 × 42 × 3, which describes the width, height, and number of layers of the input image, respectively. A batch size of 256, learning rate of 0.0001, and RMSprop optimizer (Vijendra Babu et al., 2020; Xu et al., 2021; Elshamy et al., 2023) were used for training the ResNet model for 200 epochs.
@@ 4 TSDCD evaluation
To compare the TSDCD results with those obtained in our prior investigations in WTM1, we generated a TSDCD with an $F_{low}$ of 30 Hz only with all 5 $MSNR$ ranges. This dataset consists of 4000 sources, created using identical parameters to the testing dataset in WTM1. In Figure 5, the top panel displays the accuracy of the detection acquired for the TSDCD, alongside the accuracy obtained by SSDD in WTM1. The plot clearly demonstrates a significant improvement across all five $MSNR$ ranges, particularly at lower $MSNR$ values. The detection accuracy has improved from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$, and $98.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and greater than 8, respectively. This shows a significant improvement of approximately $30.83\%$, $39.00\%$, $17.75\%$, $5.81\%$ and $1.52\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and greater than 8, respectively. It is essential to note that achieving $100\%$ accuracy for sources with $MSNR$ greater than 8 means successful detection of all sources from our samples. However, it does not guarantee that all sources within this specific range will be detected at all times. This applies to all other $MSNR$ ranges.
@@
5 Evaluation on ET’s continuous data
As previously done in WTM1, three sub-detectors combined mock data were generated to assess the efficiency of utilizing the TSDCD for near-real-time detection and to compare it with previously obtained results from SSDD. Five hours and 50 seconds of time-series data, with no overlapping sources and only BBH injections, encompassing all $MSNR$ ranges, were generated for each $F_{low}$ (5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz) for each detector. At every five-second interval in this time-series data, a signal is injected at a random position, resulting in a total of 3,610 injected sources, with each $MSNR$ range containing 722 sources. The total duration of the data for each detector is 25 hours, 4 minutes, and 16 seconds. To compare with our previously obtained results, we utilize the data with an $F_{low}$ of 30 Hz.
Figure 8: Inferencing on ET’s continuous data
@@ 5 Evaluation on ET’s continuous data
Figure 9: Top: A comparison of FPR values between SSDD (from WTM1) and TSDCD for sources with $F_{low}$ set at 30 Hz. Bottom: FPR values exclusively from TSDCD for sources with $F_{low}$ at 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz.
@@
FPR was computed for each $F_{low}$ for the entire mock data. Figure 9 depicts FPR as a function of $MSNR$, where in the top panel, TSDCD results are compared to SSDD results from WTM1. The general trend is consistent with the previous results in WTM1, where FPR decreases as $MSNR$ increases. As shown in the figure, FPR values have significantly improved, especially for sources with lower $MSNR$, transitioning from 0.712, 0.535, 0.050, 0.014, and 0.001 to 0.526, 0.296, 0.006, 0.004, and 0.0 for $MSNR$ ranges 4-5, 5-6, 6-7, 7-8, and >8, respectively. In the bottom panel, FPR values are depicted for each $F_{low}$. The overall performance is nearly identical across different $F_{low}$ settings, except for the case of $F_{low}$ at 5 Hz. This discrepancy aligns with the accuracy results presented earlier.
@@
Similarly, in terms of chirp mass $M$, the right panel in Figure 11 demonstrates a significant decrease in the total number of undetected sources in TSDCD, particularly among sources with chirp masses $M$ in the range of 15-20 $M_{\odot}$. In TSDCD alone, count of undetected sources of different $F_{low}$ varies among all chirp mass $M$ values. For $F_{low}$ at 10 Hz, the minimum count observed is 240, 98, 46,55 and 34 in the chirp mass $M$ ranges of 15-20, 20-25, 25-30,30-35 and 35-40 $M_{\odot}$, respectively. For the chirp mass $M$ ranges 40-45 and 45-50, the minimum count observed are 14 and 7 for $F_{low}$ of 20 Hz and 15 Hz, respectively. Regarding the $D_{L}$, the right panel in Figure 12 shows a substantial reduction in the count of undetected sources in TSDCD, especially for sources with $D_{L}$ in the range of 10-30 Gpc. In the left panel, no significant variation is observed across different $F_{low}$ settings. In summary, both SSDD and TSDCD exhibit a higher detection rate for sources with higher $MSNR$, chirp mass, and shorter $D_{L}$, as indicated by smaller FPR values. Notably, TSDCD demonstrates a significant improvement over SSDD. It’s worth noting that the same source may exhibit different $MSNR$ values on each sub-detector, yet TSDCD consistently outperforms SSDD across these variations.
@@
In terms of computational efficiency, we concurrently processed a total of 25 hours, four minutes, and 16 seconds of data from each detector in parallel. Utilizing the same configuration as in WTM1 (Core i7 MacBook Pro with 16 GB memory, 2667 MHz DDR4, and 2.6 GHz processor), the entire data scan was completed in 15 minutes, averaging 1.9 minutes for each hour. The processing time has been significantly reduced compared with that achieved on SSDD of 4.7 minutes for each hour. The key difference lies in simultaneously reading data from all three sub-detectors. Hence, utilizing TSDCD is suitable for near-real-time detection and could be further enhanced with a more powerful setup.
6 ET-MDC1: Einstein Telescope mock Data Challenge
A Mock Data Challenge for the Einstein Gravitational-Wave Telescope by Regimbau et al. (2012) was first released in 2012 and updated in early 2024 to contain 1.3 Terabytes of data for ET only, in addition to Cosmic Explorer. The recent release, called ET-MDC1, contains a continuous GW signal plus noise and noise only of about one month (30.8 days) split into 1300 segments of 2048 seconds, sampled at 8192 Hz. Using ET V-shaped detectors E1, E2, E3, GW signal and noise (colored Gaussian noise with no noise artifacts in this version) were simulated in addition to the null stream E0. The data contains the parameters of the injected sources, such as optimal SNR, component masses (M1 and M2) and $D_{L}$. The GW signal contains 59,540 BNS, 6,578 BBH, and 1,977 BHNS events, with optimal SNR ranges between 0.13 and 586.12, and For BNS, BBH and BHNS, IMRPhenomPNRv2 (which contains tidal effects), IMRPhenomXPHM and IMRPhenomXPHM approximant were used respectively. In ET-MDC1, BBH has a wide range of optimal SNR, spanning from 0.8 to 586. The component masses (M1 and M2) vary significantly, with minimum values of 7 and 6 $M_{\odot}$ respectively, and maximum values reaching 793 and 617 $M_{\odot}$. The $D_{L}$ of these BBH systems ranges from 0.5 Gpc to 154.37 Gpc.
The great challenge for our ResNet model to be evaluated on ET-MDC1 is that it was only trained on BBH mergers and with no overlapping sources. In addition to that, the waveforms were not continuous in the training dataset, and BNS has a long inspiral phase that can last for days, which the model has not seen during training. We evaluated ResNet on the entire ET-MDC1 E1, E2, and E3 injected data (signal plus noise) to assess its performance. Additionally, we tested the model on one week of noise-only data and null stream data to check for FPR with three different thresholds of 0.3 and 0.1 (equivalent to $50\%$, $70\%$ and $90\%$ confidence, respectively). As illustrated in Table 4, When accepting detection with at least $90\%$ confidence no FP was recorded. Only one false detection was observed using 0.3 threshold and 10 when using 0.5.
Figure 13: Optimal SNR (left) and $D_{L}$ (right) of detected and undetected BBH sources from ET-MDC1.
Figure 14: Chirp mass $M$ of detected and undetected BBH sources from ET-MDC1.
Threshold Null Noise
0.5 0 10
0.3 0 1
0.1 0 0
Table 4: FPR from one week of noise and null data only.
As shown in Figure 13 and Figure 14, ResNet successfully detected 5,566 BBH mergers out of a total of 6,578. Seventy-five percent of these sources have an average optimal SNR of 38.3, an average $D_{L}$ of 32 Gpc, and an average chirp mass of 96 $M_{\odot}$. The detected BBH sources have a minimum and maximum $D_{L}$ of 0.5 Gpc and 148.95 Gpc, respectively. The minimum and maximum optimal SNR are 1.2 and 586, respectively, and the minimum and maximum chirp mass are 6 and 596 $M_{\odot}$, respectively.
For undetected BBH sources, the optimal SNR ranges from 0.8 to 51.7, with seventy-five percent having an average optimal SNR of 10.7. The $D_{L}$ ranges between 2.4 Gpc and 154.4 Gpc. The minimum and maximum chirp mass of the undetected sources are 7.8 $M_{\odot}$ and 484.6 $M_{\odot}$, respectively.
These results demonstrate the great performance of our model. The performance can be significantly improved by considering parameters outside the current range for BBH masses and distances. Additionally, incorporating samples of overlapping sources into the training dataset will positively impact the results.
Figure 15: Optimal SNR of detected and undetected BNS (right) and BHNS (left) sources from ET-MDC1.
Although ResNet was not trained on BNS and BHNS mergers, the model was able to detect 11,477 BNS mergers and 323 BHNS mergers. The optimal SNR of the detected and undetected BNS and BHNS sources are shown in the right and left panel of Figure 15. BNS and BHNS detected sources have optimal SNR ranges from 0.2 to 383 and from 1 to 50, respectively.
The observation that the ResNet model is capable of detecting BNS and BHNS mergers, despite being trained only on BBH mergers, is due to the fact that inspiral GW signals share a common functional structure consisting of three phases: inspiral, merger, and ringdown. However, the exact form of the signal depends on several physical parameters such as the masses of the binary objects, their spins, the orbital eccentricity, and the distance to the source. When using matched filtering, a template generated with a specific set of parameters may perform poorly for detecting a signal with a different set of parameters, leading to a reduction in cross-correlation between the template and the observed signal (Macleod et al., 2016). In contrast to matched filtering, deep learning models have the capacity to generalise beyond their training data. This characteristic allows CNNs to capture shared features across different types of compact binary systems, including BBH, BNS, and BHNS. All of these systems produce signals that include inspiral, merger, and ringdown phases, making the generalisation ideal candidates for CNN-based classification. When training our model, we specifically focused on the merger phase, which is common across all compact binary systems. Our model was trained on thousands of merger waveforms with component masses ranging from 10 to 56 solar masses. This training allowed the network to learn features that generalise well across systems, irrespective of the specific mass or spin configurations. However, the performance can be significantly improved if this model is trained on diverse and well representatives samples of all three different systems, in addition to overlapping and non-overlapping sources.
7 PyMerger
PyMerger (Alhassan et al., 2024) is a Python tool for detecting BBH mergers from ET, built based on the trained ResNet model described above. PyMerger is our first step towards developing a full AI-based pipeline for detection and, soon, parameter estimation of BBH from ET. We developed and shared the first version to open-source the trained model and allow results reproducibility.
Figure 16: A flowchart outlines the high-level process of PyMerger, starting from argument parsing to data processing, model prediction, and result logging.
The current version handles only gravitational wave frame file format (.gwf), with the intention to support HDF5 in the upcoming version. As shown in Figure 16, the software takes seven arguments: 1) Directory containing the input .gwf files. 2) Directory to store the results. 3) Sampling-rate: Sampling rate of the input data (either 8192 or 4096). Default is 8192. 4) Number of data segments to be processed for each detector (i.e., number of .gwf files to be processed for each sub-detector). Default is 1. 5) List of the three channels names to be processed. Default is [’E1:STRAIN’, ’E2:STRAIN’, ’E3:STRAIN’]. 6) Threshold value for merger detection. Default is 0.1 (accepting detection with at least $90\%$ confidence). The software processes each file by sliding a window (window size based on the sampling rate) to scan the data and look for BBH mergers. The output is a text file (.txt) with three columns: 1) Starting GPS time of the sliding window. 2) End GPS time of the sliding window. 3) Detection probability of all the detected mergers.
Regarding computing efficiency, PyMerger scans one hour of TSDCD in 1.9 minutes. This can be significantly improved, as deep learning models can be substantially accelerated based on the hardware used. For instance, using the NVIDIA TensorRT-optimized AI ensemble, Chaturvedi et al. (2022) processed one month of advanced LIGO data in 50 seconds. Future work will focus on incorporating BNS and BHNS events along with glitchy noise.
The software is publicly available at https://github.com/wathela/PyMerger. Researchers in the field are encouraged to make use of the software and kindly provide us with their feedback, comments, and suggestions.
@@ 8 Conclusion
In this study, we conducted a comparative analysis of the efficiency in detecting BBH sources using data from all three sub-detectors of ET simultaneously versus using data from a single sub-detector. The comparison was specifically focused on cases with $F_{\text{low}}$ set to 30 Hz. Furthermore, we explored TSDCD’s performance across five different $F_{\text{low}}$ settings (5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz), employing the same $MSNR$ ranges as in our previous study (4-5, 5-6, 6-7, 7-8, and >8). Utilizing the ResNet model, which exhibited superior performance in our previous study, TSDCD demonstrated a significant enhancement in detection accuracy compared to SSDD. The accuracy improved from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$, and $98.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ ranges of 4-5, 5-6, 6-7, 7-8, and >8, respectively. The results indicate a substantial accuracy improvement for lower $MSNR$ ranges: 4-5, 5-6, and 6-7, with gains of $18.5\%$, $24.5\%$, and $13\%$ respectively. Additionally, there is a significant improvement of $5.5\%$ and $1.5\%$ for higher $MSNR$ ranges: 7-8 and >8.
For more rigorous evaluation, ResNet model was evaluated on ET-MDC1 dataset, where the model demonstrated strong performance in detecting BBH mergers, identifying 5,566 out of 6,578 BBH events, with optimal SNR starting from 1.2, and a minimum and maximum $D_{L}$ of 0.5 Gpc and 148.95 Gpc, respectively. Despite being trained only on BBH mergers without overlapping sources, the model achieved high detection rates. However, to further enhance its performance, it is essential to include a broader range of parameters for BBH masses and distances, and incorporate overlapping sources in the training data. Notably, even though the model was not trained on BNS and BHNS mergers, it successfully detected 11,477 BNS and 323 BHNS mergers, with optimal SNR starting from 0.2 and 1, respectively. This indicates the model’s potential for broader applicability. Future work will focus on improving detection rates by including both overlapping and non-overlapping BNS and BHNS events in the training process. Built on ResNet, PyMerger is a Python tool designed for detecting BBH mergers from ET data. PyMerger operates without the need for data whitening or band-passing and can scan one hour of TSDCD in 1.9 minutes on an average laptop. The current version supports GW frame format (.gwf) files.
@@
This research was funded in part by National Science Centre, Poland (UMO-2023/49/N/ST9/01287). Additionally, This work has been supported in part by the European Union’s Horizon 2020 research and innovation programme under grant agreement No 952480 (DarkWave project), and from the International Research Agenda Programme AstroCeNT (MAB/2018/7) funded by the Foundation for Polish Science (FNP) from the European Regional Development Fund.
Appendix A ResNet: Deep Residual Neural Networks
Convolutional Neural Networks (CNNs) Fukushima (1980) are deep learning models designed to process data with a 2D structure, like images or spectrograms, for tasks such as pattern recognition and object detection O’Shea & Nash (2015). A CNN typically consists of convolutional layers that extract features using filters applied through convolution operations Goodfellow et al. (2016), followed by activation functions (e.g., ReLU, $tanh$, $softmax$) Haykin (1998); Bridle (1990). Pooling layers reduce the spatial dimensions by aggregating local regions, and fully connected layers act as classifiers, similar to traditional Artificial Neural Networks (ANNs) Agarap (2017). In essence, CNNs perform non-linear mappings from input images to class scores. CNNs model suffer from vanishing gradient problem (also called degradation problem), where increasing the depth of the network leads to higher training error. When deep neural networks are trained, an increase in network depth often leads to a saturation and subsequent degradation in performance. The degradation is not caused by overfitting but by difficulties in optimization. This is counterintuitive because deeper models are expected to, theoretically, learn better representations.
Deep Residual Neural Networks (ResNets) were introduced by He et al. (2015) as a solution to the degradation problem in deep learning models. ResNets solve this problem by introducing shortcut connections that bypass certain layers, allowing for easier optimization of deep models. ResNet introduces the concept of residual learning.The core component of ResNet is the residual block, shown in Figure 17, where the network learns a residual function instead of directly learning the desired mapping. Let $X$ represent the input to a residual block and $\mathcal{F}(X)$ be the desired transformation. In a traditional network, the output of a layer would be $H(X)=\mathcal{F}(X)$. However, in ResNet, the network learns a residual function $\mathcal{R}(X)$ defined as:
$\mathcal{R}(X)=\mathcal{F}(X)-X,$ (A1)
which can be rearranged as:
$\mathcal{F}(X)=\mathcal{R}(X)+X.$ (A2)
This residual function $\mathcal{R}(X)$ is learned by a series of layers, and the output of the residual block is:
$Y=X+\mathcal{R}(X).$ (A3)
The addition of the input $X$ to the learned residual mapping allows the network to bypass transformations, helping to preserve gradients during backpropagation.
Figure 17: Residual learning: a residual block.
By stacking multiple residual blocks, ResNet enables the training of very deep networks without the performance degradation seen in traditional deep architectures. The general form for the output of the $l$-th residual block in a deep ResNet can be written as:
$Y^{[l+1]}=Y^{[l]}+\mathcal{F}(Y^{[l]},W^{[l]}),$ (A4)
where $Y^{[l]}$ is the input to the $l$-th block, $\mathcal{F}(Y^{[l]},W^{[l]})$ is the residual function learned by the $l$-th block, and $W^{[l]}$ are the trainable weights of the block. For deeper networks like ResNet-50 or ResNet-101, the bottleneck design is used to reduce the number of parameters. In this design, three layers are used instead of two. These layers use a combination of three convolutions: 1) A 1x1 convolution to reduce the dimensionality 2) A 3x3 convolution for feature extraction 3) A final 1x1 convolution to restore the original dimensionality. The residual function for a bottleneck block is:
$\mathcal{R}(X)=W_{3}\cdot(W_{2}\cdot(W_{1}\cdot X)),$ (A5)
where $W_{1}$, $W_{2}$, and $W_{3}$ are the weights for the three convolutional layers. The bottleneck design reduces computational complexity by reducing the dimensionality of the feature maps before applying expensive convolutions. When the dimensions of the input $X$ and the output $\mathcal{F}(X)$ do not match, ResNet employs a projection shortcut to match dimensions using a linear projection:
$Y=W_{s}\cdot X+\mathcal{F}(X),$ (A6)
where $W_{s}$ is the projection matrix (typically a 1x1 convolution) used to adjust the dimensions of $X$ before adding it to the residual function. For the training, ResNet uses standard backpropagation (Munro, 2010), with the total loss $\mathcal{L}$ minimized using stochastic gradient descent (SGD). For classification tasks, the loss function is typically cross-entropy:
$\mathcal{L}=-\sum_{i}y_{i}\log(\hat{y}_{i}),$ (A7)
where $y_{i}$ is the true label and $\hat{y}_{i}$ is the predicted probability for class $i$. The introduction of residual connections significantly improved the performance of deep neural networks. ResNet achieved state-of-the-art results in the ImageNet competition (Russakovsky et al., 2015), and has been widely adopted in various applications, including image classification, object detection, and segmentation.